OpenStudy (anonymous):

Is this a prime? x^2+14x+5

OpenStudy (mertsj):

Are there factors of 5 whose sum is 14?

OpenStudy (anonymous):

well i noticed it doesnt factor very well. (x+7)^2 - 7^2 + 5 = 0 (x+7)^2 = 44 x+7 = √44 x= -7±√44 So irrational "factors" would be: (x+7+√44)(x+7-√44) ..

OpenStudy (mertsj):

So normally, it would be classified as prime.

OpenStudy (anonymous):

it is good that you can see that, if you ever start playing around with something called Sophie Germain Identity, but I'd bet you professor is looking for you to call it prime.

OpenStudy (anonymous):

i mean, just by definition, a prime number must (among other things) be a positive integer, right?

OpenStudy (anonymous):

so maybe that rules out irrational "factors"

OpenStudy (anonymous):

I know, I know I thought so but i just wanted to make sure lol before I input the answer.

OpenStudy (precal):